ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 165, 2008

Abrams, G., G. Aranda Pino and M. Siles Molina Locally finite Leavitt path algebras

329

Ammar, M. and N. Poncin Formal Poisson cohomology of twisted r-matrix induced structures

381

Aranda Pino, G. See G. Abrams

Badea, C. and S. Grivaux Faber-hypercyclic operators  43

Boyarchenko, M. and M. Sabitova The orbit method for profinite groups and a p-adic analogue of Brown’s theorem

67

Choi, SoYoung Linear relations and congruences for the coefficients of Drinfeld modular forms

93

Downarowicz, T. Faces of simplexes of invariant measures  189

Grivaux, S. See C. Badea

Hundley, J. Spin L-functions for GSO_{10} and GSO_{12}  103

Mahalanobis, A. The Diffie-Hellman key exchange protocol and nonabelian nilpotent groups

161

Marzantowicz, W. and F. Przytycki Entropy Conjecture for continuous maps of nilmanifolds

349

Miana, P. J. Algebra homomorphisms from cosine convolution algebras

253

van Mill, J. Homogeneous spaces and transitive actions by Polish groups

133

Naveh, Y. Tight upper bounds on the number of invariant components on translation surfaces

211

Pinchover, Y., G. Wolansky and D. Zelig Spectral properties of Schrödinger operators defined on N-dimensional infinite trees

281

Poncin, N. See M. Ammar

Przytycki, F. See W. Marzantowicz

Roychowdhury, M. K. and D. J. Rudolph Any two irreducible Markov chains of equal entropy are finitarily Kakutani equivalent

29

Rudolph, D. J. See M. K. Roychowdhury

Sabitova, M. See M. Boyarchenko

Siles Molina, M. See G. Abrams Singh, A. and M. Thakur Reality properties of conjugacy classes in

algebraic groups

1

Thakur, M. See A. Singh Wolansky, G. See Y. Pinchover Zelig, D. See Y. Pinchover

Zhang, Gaofei On the non-escaping set of e^{2πiθ} sin(z)  233